# Generic (fractional) quantum anomalous Hall crystals from interaction-driven band folding

### v1

@@ Abstract
We then turn to the fractional fillings of the miniband. Apart from confirming the previous possible states by ED, we find a series of new FQAHC states and we will take two as examples for illustration with more details in the SI sup . We show the spectral flow of a 30-site torus at $\nu=11/15$ in Fig.2 (c) and find there are 15 degenerate ground states and the Hall conductivity (by calculating the many-body Chern number) of each ground state is $\sigma_{\mathrm{H}}=-\nu^{\ast}=-1/5$. The topological degeneracy is $5$ while each of them is tripled due to the CDW order, and these ground states flow back to themselves after inserting $10\pi$ fluxes, which further suggests FQAHC state is a Laughlin state with the coexisting CDW order. As mentioned earlier, the robustness of FQAHC states should be examined by larger-scale simulations and we show the $N_{y}=6$ iDMRG results at $\nu=7/9$ in Fig.2 (d-f). After adiabatically inserting $6\pi$ fluxes, one electron is pumped and the entanglement spectrum (ES) is shifted by one charge Li and Haldane (2008); Alexandradinata et al. (2011), supporting the fractional Hall conductivity $\sigma_{\mathrm{H}}=-\nu^{\ast}=-1/3$. Further, the momentum-resolved ES shows the characteristic edge-mode counting $\{1,1,2,3,5,...\}$, in agreement with the edge conformal field theory Li and Haldane (2008); Cincio and Vidal (2013); Regnault (2015), which further provides the smoking gun that the topological order in the FQAHC state could coexist with the CDW order via interaction-driven band folding. Besides, although the dispersion and geometry of the miniband is worse with only $V_{1}$ as discussed, we find it still possible to host some FQAHC states when fractionally filled, in contrast with the previous ED study Kourtis and Daghofer (2014); sup .
@@
Thermodynamics of the FQAHC state.— Although the thermodynamic properties of symmetric FQAH states have been studied Lu et al. (2024c, d), finite-temperature knowledge of the FQAHC state with intertwined orders is still elusive, so we take the $\sigma_{\mathrm{H}}=-1/3$ FQAHC state as an illustration. The $\bar{n}-\mu$ plateaus at $\nu=2\bar{n}=7/9\ (\nu^{\ast}=1/3)$ are shown in Fig.3(a), supporting the incompressibility at low temperatures. The compressibility and specific heat as functions of temperature are shown in Fig.3 (b,c), respectively. We find that the onset temperature of the incompressibility of the FQAHC phase is at $T^{\ast}\sim 0.14$, around which there is a small peak in specific heat and both quantities show exponential decay upon further decreasing temperature. In addition, the $\bar{n}-\mu$ plateau shrinks with increasing temperatures and is about to vanish above $T^{\ast}$, further supporting $T^{\ast}$ as the onset of incompressibility. As the FQAHC phase has a coexisting CDW order, we further focus on its formation. The CDW order parameter $\rho(\frac{2\pi}{3},0)$ with $\rho(\mathbf{k})=\frac{1}{N_{i}}\sum_{i}e^{i\mathbf{k}\mathbf{r}_{i}}(\langle n_{i}^{\mathrm{A}}\rangle-\bar{n})$ (take the A sublattice for example and $N_{i}$ is the number of counted sites), and the structure factor $S(\frac{2\pi}{3},0)$ (referring to such CDW fluctuations), are shown in Fig.3 (d). The CDW fluctuations reach their peak at $T_{\mathrm{CDW}}\sim 1$ (where the CDW order forms), and this CDW transition temperature is higher than the onset temperature of the incompressible FQAHC phase, suggesting a compressible CDW phase at intermediate temperature $T^{\ast}<T<T_{\mathrm{CDW}}$, and the symmetric phase exists at higher temperature $T>T_{\mathrm{CDW}}$. We also show the structure factors in the original BZ in Fig.3(e,f). Around $T_{\mathrm{CDW}}$, there exist sharp peaks at ±$\pm(\frac{2\pi}{3},0)$. However, at $T=0.07<T^{\ast}$, as the CDW order is already well established, we observe that the momenta of the charge fluctuations are no longer at the CDW vectors but around the boundaries of the BZ, which might be related to the low-energy magnetoroton mode (intrinsic collective excitations in the FQH and FQAH states) Girvin et al. (1986); Lu et al. (2024c, d); Zaklama et al. (2024); Long et al. (2025).
@@
Discussions.— In conclusion, we have illustrated a generic mechanism to realize a series of (F)QAHCs from interaction-driven band folding and renormalization, which offers an intriguing combination of the quantum-Hall and Mott physics. For the FQAHC states here, we find them qualitatively different from the previously found FQAH+CDW state (realized from roton softening in the symmetric FQAH state) Lu et al. (2025b). The Hall conductivity of the latter is equal to the filling of the Chern band (with $C=1$) and that of the symmetric FQAH state, while the former states have $\sigma_{\mathrm{H}}=\nu^{\ast}\neq\nu$, which might suggest that only part of the electrons in the FQAHC states (from band folding) contribute to the transport while the other part mainly form the CDW. However, for both situations, there is a similarity that the CDW formation temperature is higher than the temperature scale of the onset of incompressibility, such that there are compressible CDW states at intermediate temperature range Lu et al. (2024b).
@@
More importantly, the different roton-driven and band-folding mechanisms together help with the more systematic understanding of the FQAH+CDW states. Interestingly, we note that these two origins of the FQAH+CDW states are very similar to two mechanisms of realizing supersolids : roton softening in the superfluid Chomaz et al. (2019), and doping the solids with the condensation of the dopants Chen et al. (2008). Such analogy of the FQAH+CDW with supersolid physics, further deepens our understanding of these exotic states of quantum matter beyond traditional liquid and solid phases.
In this work, we have also found that the competing interactions could flatten the energy dispersion and smoothen the quantum geometry of the CDW-driven topological miniband, possibly due to the effectively introduced longer-range couplings. As possible Hall crystals at fractional fillings due to band-folding have been experimentally reported Su et al. (2025), it would be interesting to detect and study how to improve the quantum geometry of the folded bands, in order to realize more exotic FQAHC states upon doping (similar to what we showed here). Apart from the moiré materials, it would also be interesting to explore such band-folding schemes for topological models in cold-atom systems Léonard et al. (2023).
Although we mainly focus on large $V_{1}$ interactions for clear illustration and this CDW at $\nu=2/3$ and different interaction strengths is adiabatically connected, it woule be still interesting for future work to numerically study the weak limit of the CDW order for this scheme of (F)QAHC states. Further, as we focus on a system with spinless fermions and discrete symmetry, it would be interesting to explore the elusive ferromagnetic FQAHC state and the FQAHC state with continuous translation-symmetry breaking, which might render more complex yet richer finite-temperature properties.
We also notice the recent discovery of an extended fractional Chern insulator in twisted bilayer graphenes (Hofstadter regime), possibly due to the doped quasiparticles forming crystals Finney et al. (2025). It would be interesting to explore in microscopic models if the doped quasiparticles from the FQAH states could form crystals and coexist with the original FQAH states, and compare such states to the FQAHCs in this work.
@@ Acknowledgements.
Acknowledgments —  We thank Maria Daghofer, B. Andrei Bernevig, Nicolas Regnault, Jie Wang, and Yang Liu for helpful discussions. HYL thank Shou-Shu Gong and Xiao-Tian Zhang for discussing the DMRG simulations and Wenqi Yang for discussing the mean-field calculations. The TeNPy package is used for the infinite DMRG simulations Hauschild and Pollmann (2018) and the QSpace package is used for the thermodynamic simulations Weichselbaum (2012). HYL and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper. H.Q. Wu acknowledge the support from National Natural Science Foundation of China (No. 12474248), GuangDong Basic and Applied Basic Research Foundation (No. 2023B1515120013), Guangdong Fundamental Research Center for Magnetoelectric Physics (Grant No.2024B0303390001), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008). WY acknowledges support by the National Natural Science Foundation of China (No. 12425406), Research Grant Council of Hong Kong (AoE/P-701/20, HKU SRFS21227S05), and New Cornerstone Science Foundation.

### v3

@@ Abstract
Therefore, we start from the simpler case with $V_{1}=10,\ V_{2}=V_{3}=2$. Apart from confirming the $|\sigma_{\mathrm{H}}|=\frac{2}{5},\ \frac{3}{5}$ FQAHCs (found by ED), we find a series of new FQAHC states at fractional fillings of this C=1C=-1 mini-band and (more details in the SI sup ). We show the spectral flow of a 30-site torus at $\nu=11/15$ in Fig.2 (c) and find there are 15 degenerate ground states and the Hall conductivity (by calculating the many-body Chern number) of each ground state is $\sigma_{\mathrm{H}}=-\nu^{\ast}=-1/5$. The topological degeneracy is $5$ while each of them is tripled due to the CDW order, and these ground states flow back to themselves after inserting $10\pi$ fluxes, which further suggests FQAHC state is a Laughlin state with the coexisting CDW order. However, according to the band-folding scenario, the finite-size effect for ED might be more severe than that for symmetric FQAH states. For example, in the 30-site torus with 15 unit cells, there are only 5 momenta in the C=1C=-1 mini-band. Thus, the robustness of FQAHC states should be examined by larger-scale simulations and we show the $N_{y}=6$ iDMRG results at $\nu=7/9$ and $V_{1}=10,\ V_{2}=V_{3}=2$ in Fig.2 (d-f). After adiabatically inserting $6\pi$ fluxes, one electron is pumped and the entanglement spectrum (ES) is shifted by one charge Li and Haldane (2008); Alexandradinata et al. (2011), supporting the fractional Hall conductivity $\sigma_{\mathrm{H}}=-\nu^{\ast}=-1/3$. Further, the momentum-resolved ES shows the characteristic edge-mode counting $\{1,1,2,3,5,$...\}, in agreement with the edge conformal field theory Li and Haldane (2008); Cincio and Vidal (2013); Regnault (2015), which further provides the smoking gun that the topological order in the FQAHC state could coexist with the CDW order via interaction-driven band folding.
The next question is whether there exists FQAHCs with only $V_{1}$ as the mini-band is topological in principle. At $V_{1}=10$, we only find evidence of the $\sigma_{\mathrm{H}}=-1/3$ FQAHC at $\nu=7/9$, and the charge pumping result is shown in Fig. 2(d), while the curve is less straight than that with $V_{2}=V_{3}=2$, which might suggest the less uniform many-body Berry curvature Sheng et al. (2006); Lu et al. (2022). According to the further Hartree-Fock analysis of the C=1C=-1 mini-band, compared to the case at $V_{2}=V_{3}=0$, the dispersion is flatter and the quantum geometry is more uniform at $V_{2}=V_{3}=2$ sup , which might help explain why the other FQAHCs are more easily realized there.
Figure 4: Bosonic FQAHC at $\nu=5/6$ at $V_{1}=10,\ V_{2}=V_{3}=0$. (a) The momentum-resolved ES of the $\sigma_{\mathrm{H}}=-1/2$ FQAHC from iDMRG simulation of the $N_{y}=4$ infinite cylinder. (b) The charge pumping result of the $\sigma_{\mathrm{H}}=-1/2$ FQAHC.
@@
Thermodynamics of the FQAHC state.— Unlike the thermodynamics of symmetric FQAH states Lu et al. (2024c, d), the finite-temperature knowledge of the FQAHC state with intertwined orders is still elusive, so we showcase the finite temperature properties of the $\sigma_{\mathrm{H}}=-1/3$ FQAHC state at $V_{1}=10,\ V_{2}=V_{3}=2$. The $\bar{n}-\mu$ plateaus at $\nu=2\bar{n}=7/9\ (\nu^{\ast}=1/3)$ are shown in Fig.3(a), supporting the incompressibility at low temperatures. The compressibility and specific heat as functions of temperature are shown in Fig.3 (b,c), respectively. We find that the onset temperature of the incompressibility of the FQAHC phase is at $T^{\ast}\sim 0.14$, around which there is a small peak in specific heat and both quantities show exponential decay upon further decreasing temperature. In addition, the $\bar{n}-\mu$ plateau shrinks with increasing temperatures and is about to vanish above $T^{\ast}$. We then further focus on the formation of CDW. The CDW order parameter $\rho(\frac{2\pi}{3},0)$ with $\rho(\mathbf{k})=\frac{1}{N_{i}}\sum_{i}e^{i\mathbf{k}\mathbf{r}_{i}}(\langle n_{i}^{\mathrm{A}}\rangle-\bar{n})$ (sublattice $A$ for example and $N_{i}$ is the number of counted sites), and the structure factor $S(\frac{2\pi}{3},0)$ (referring to such CDW fluctuations) with $S(\mathbf{q})=\sum_{ij}e^{-i\mathbf{q}(\mathbf{r_{j}}-\mathbf{r_{i}})}(\langle n_{i}^{\mathrm{A}}n_{j}^{\mathrm{A}}\rangle-\langle n_{i}^{\mathrm{A}}\rangle\langle n_{j}^{\mathrm{A}}\rangle)$, are shown in Fig.3 (d). The CDW fluctuations reach their peak at $T_{\mathrm{CDW}}\sim 1$ (where the CDW order forms), and this CDW transition temperature is higher than the onset temperature of the incompressible FQAHC phase, suggesting the compressible CDW phase at intermediate temperature $T^{\ast}<T<T_{\mathrm{CDW}}$, and the symmetric phase at higher temperature $T>T_{\mathrm{CDW}}$. Despite the CDW order, the low-energy magnetoroton mode (intrinsic collective excitations in the FQ(A)H states Girvin et al. (1986); Lu et al. (2024c, d); Zaklama et al. (2024); Long et al. (2025).) at the different momentum from the CDW vector could still be observed sup .
Bosonic FQAHC state.— Furthermore, we show that this generic band-folding scheme for FQAHCs could be applied to bosonic systems. Here, we consider this two-band model filled with hard-core bosons (no more than 1 boson per site) by replacing the fermionic creation/annihilation operators in Eq. (1) with bosonic ones. The same trivial CDW at $\nu=2/3$ is observed and we show the iDMRG results at $\nu=5/6$ ($\nu^{\ast}=1/2$) in Fig.4. The characteristic counting of edge modes is demonstrated, and after adiabatically inserting $4\pi$ flux, exactly one boson is pumped from right to the left side, supporting the bosonic $\sigma_{\mathrm{H}}=-1/2$ FQAHC state.
@@
Discussions.— In this work, we demonstrate a generic mechanism to realize a series of (F)QAHCs from interaction-driven band folding and renormalization effects, which offers an intriguing combination of the gems of quantum-Hall and Mott physics. For the FQAHC states here, we find them qualitatively different from the previously found FQAH+CDW state (realized from roton softening in the symmetric FQAH state) Lu et al. (2025b). The Hall conductivity of the latter is equal to the filling of the Chern band (with $C=1$) and that of the symmetric FQAH state, while the former states have $\sigma_{\mathrm{H}}=\nu^{\ast}\neq\nu$, which might suggest that only part of the electrons in the FQAHC states (from band folding) contribute to the transport while the other part mainly form the CDW.
@@
More importantly, the different roton-driven and band-folding mechanisms together help with the more systematic understanding of the FQAH+CDW states. We note that these two origins of the FQAH+CDW states are very similar to two mechanisms of realizing supersolids: roton softening in the superfluid Chomaz et al. (2019), and doping the solids with the condensation of the dopants Chen et al. (2008). Such analogy of the FQAH+CDW with supersolid physics, further deepens our understanding of these exotic states of quantum matter beyond traditional liquid and solid phases.
While we realize FQAHCs by doping the topological mini-band above the gap of the trivial commensurate CDW, it is interesting to study whether they can be realized by dope the integer QAHC states at commensurate fillings. Besides, our work suggests the importance of quantum geometry of CDW-driven topological mini-band, which is also mentioned while not explored in recent QAHC experiments Su et al. (2025). To improve the quantum geometry, while we use competing longer-range interactions here, the method might vary in realistic experiments, such as the weak magnetic field for native Chern bands Xie et al. (2021).
The compressible CDW phase at the intermediate temperature of the FQAHC is also interesting as is might exhibit anisotropic transport properties depending on the specific CDW order, which manifest in the longitudinal resistivity. Reversely, such a finite-temperature CDW metal can be viewed as a precursor of the lower-temperature Hall crystal physics in FQAH systems. Although the metallic CDW phase would not have quantized Hall resistivity, whether there exist nontrivial features of the transverse resistivity at the finite-temperature region of the FQAHC state is interesting for future studies.
Moreover, we have extended and broadened the platform of Hall crystals to bosonic systems. Considering the progress made in realizing bosonic FQH states in cold-atom systems, we believe our work demonstrates a possible scheme to realize FQAHC states in such systems Léonard et al. (2023); Wang et al. (2024b).
@@ Acknowledgements.
Acknowledgments —  We thank Maria Daghofer, Kai Sun, B. Andrei Bernevig, Nicolas Regnault, Jie Wang, and Yang Liu for helpful discussions. HYL thank Shou-Shu Gong and Xiao-Tian Zhang for discussing the DMRG simulations and Wenqi Yang for discussing the mean-field calculations. The TeNPy package is used for the infinite DMRG simulations Hauschild and Pollmann (2018) and the QSpace package is used for the thermodynamic simulations Weichselbaum (2012). HYL and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as Beijing Paratera Tech Corp., Ltd par for providing HPC resources that have contributed to the research results reported within this paper. H.Q. Wu acknowledge the support from National Natural Science Foundation of China (No. 12474248), GuangDong Basic and Applied Basic Research Foundation (No. 2023B1515120013), Guangdong Fundamental Research Center for Magnetoelectric Physics (Grant No.2024B0303390001), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008). WY acknowledges support by the National Natural Science Foundation of China (No. 12425406), Research Grant Council of Hong Kong (AoE/P-701/20, HKU SRFS21227S05), and New Cornerstone Science Foundation.